OpenStudy (anonymous):

What is the 7th number in the pattern: 1, 4, 9, 16, 25, …? A. 32 B. 36 C. 39 D. 49

jimthompson5910 (jim_thompson5910):

This is the list of perfect squares. So the 7th perfect square is...?

OpenStudy (anonymous):

an=a1+(n-1)

OpenStudy (anonymous):

Huh?

jimthompson5910 (jim_thompson5910):

The 7th perfect square is 7^2 = 7*7 = 49 So the 7th number in the pattern is 49

OpenStudy (anonymous):

sorry for d at the end

OpenStudy (anonymous):

Ohhhhhh!! Thanks! (: